General Meeting Summary 10/11/17
Jump to navigation
Jump to search
DETECTORS AT JLAB (Vladimir, Hamlet, Greg)
- SHMS aerogel cherenkov detector
- Further analysis shows the need to changed PMT #6 negative before run starts
- Additional noise observed in cosmic run data may be due to bases - remove one PMT and check in dark box with LED
- Also check HV channels
- Note on thresholds: threshold during KPP (with beam) was 10mV, now have much lower threshold - may see features like reflections that would have never seen in KPP run
- Neutral Particle Spectrometer
- 160 more crystals arrived - total is 320 now
- Visual inspection of one randomly selected box from this batch shows similar trends as Box #3, i.e. about half of the crystals appear to have issues
- Crystals 5031-40 and 5061-70, and box #4 (30 total) are at CUA
- Mechanical measurements - waiting for quote from company
- Crystal database - looking into best software to record information for each crystal
- DIRC
- New postdoc expected to start in November. Responsibilities will include analysis of Gluex data (discussions with analysis coordinator ongoing) and EIC part of the DIRC (planning to spend some time at GSI and learn measurement methods)
- Good progress with preparation of BaBar DIRC bar transportation
CERN DETECTORS (Rachel)
- Funds for phase II available - procurement can be initiated
- Discussion about problems with pixel detector at CERN
- Symptom are logic board upsets
- Seems to be a design issue as one of circuits was not designed radiation hard
- Investigating possible options to address this problem (short term and longer term)
DATA ANALYSES (Salina, Andres, Richard, Vladimir, Marco)
- Discussion about Salina's DVCS analysis - compare ADC (HRS electron arm) and ARS (photon calorimeter) time windows
- TDIS - Marco will present his work in a future meeting
- E12-09-011 - making progress